OpenStudy (anonymous):

A large industrial firm uses three local motels to provide overnight accommodations for its clients. From past experience it is known that 20% of the clients are assigned rooms at the Ramada Inn, 50% at the Sheraton, and 30% at the Lakeview Motor Lodge. If the plumbing is faulty in 5% of the rooms at the Ra- mada Inn, in 4% of the rooms at the Sheraton, and in 8% of the rooms at the Lakeview Motor Lodge, what is the probability that (a) a client will be assigned a room with faulty plumbing? (b) a person with a room having faulty plumbing was assigned accommodations at the Lakeview

OpenStudy (kuoministers):

= 0.17

OpenStudy (kuoministers):

From past experience it is known that 20%of the clients are assigned rooms at Ramada Inn,50% at theSheraton, and 30% at the Lakeview Motor Lodge. and If the plumbing is faulty in5% of the rooms at Ramada Inn, in 4% of the rooms T Sheraton, andin 8% of the rooms at the Lakeview Motor Lodge, i.e, faulty 20% Ramada 5% 50% Sherata 4% 30% Lakeview 8%

OpenStudy (kropot72):

(a) Pr = probability of being assigned to the Ramada and getting a room with faulty plumbing = 0.2 * 0.05 = 0.01 Ps = probability of being assigned to the Sheraton and getting a room with faulty plumbing = 0.5 * 0.04 = 0.02 Pl = probability of being assigned to the Lakeview and getting a room with faulty plumbing = 0.3 * 0.08 = 0.024 The probability that a client will be assigned a room with faulty plumbing = Pr + Ps + Pl = 0.01 + 0.02 + 0.024 = ?